# Gravitational Lensing of Gravitational Waves from Astrophysical Sources: Theory, Detection, and Applications

### 07 May 2026 v1

@@ 2.2.2 Wave Optics
In the wave-optics regime, i.e., ${\lambda_{\rm GW}}/{\lambda_{\rm E}}\gtrsim 1$, the diffraction of the GWs becomes important, which may also lead to changes of the waveform in both amplitude and phase. Here we briefly review the basic theory of the gravitational lensing of GWs in the wave-optics regime. Figure 2 shows a schematic diagram for the lensing of a GW event by a (dark matter) minihalo in the wave-optics regime. More detailed discussions can be seen in Takahashi & Nakamura (2003).
@@
where $g^{(\rm B)}_{\mu\nu}$ is the background metric tensor. Consider a linear perturbation $h_{\mu\nu}$ in the background metric $g^{(\rm B)}_{\mu\nu}$, the metric is then
@@
where $n_{i}=0$, $0.5$, and $1$ when $x_{j}$ is the minimum, saddle, and maximum points of $t_{\rm d}(\boldsymbol{x,y})$, respectively, representing the parity of the images. This equation reduces to the geometric-optics expression in the short-wavelength limit, implying that geometric-optics is a valid approximation when diffraction can be ignored.
We note here that calculating lensed waveforms in the wave-optics regime via the amplification factor $F(f)$ via Equation (41) presents two key challenges. First, from a theoretical standpoint, Equation (41) is derived using Kirchhoff’s diffraction theory, which inherently assumes that the incident GW is obstructed by the lens and only accounts for scattered wave contribution (see also Fig. 2 for a schematic illustration). When considering that GWs propagate past BHs, fully numerical relativity simulations are necessary to obtain more robust lensed GW waveforms that incorporate the contributions from incident GWs, rather than only scattered components (e.g., He 2021; He & Wu 2022; Yin & He 2024). Second, from a numerical computational perspective, calculating $F(f)$ via Equation (41) in the wave-optics regime is both challenging and computationally expensive owing to highly oscillatory nature of the diffraction integral. To date, efficient computation of this integral is only feasible for axis-symmetric lens profiles Guo & Lu (2020).
@@ 3.1.1 Compact Binary Coalescence
Binary neutron star (BNSs) mergers, black hole-neutron star (BHNS) mergers and a significant fraction of stellar binary black holes (sBBHs) are believed to originate from the evolution of massive binary stars (hereafter, EMBS channel). One may adopt the cosmic merger rate densities directly obtained from GW observations, however, the uncertainties in these measurements are still large due to small number of detected GW events and lack of high redshift GW events. A widely adopted approach is to combine compact binary population synthesis (CBPS) model and cosmological simulation of galaxy formation and evolution (e.g., Dominik et al. 2012, 2013, 2015; de Mink & Belczynski 2015; Mapelli & Giacobbo 2018; Han et al. 2020; Chu et al. 2022; Fragos et al. 2023; Chu et al. 2022; Chen et al. 2024c, 2025b, also see recent reviews by Mandel & Broekgaarden 2022 and Lu 2025a). In principle, the cosmic merger rate density $R^{\rm S}(m_{1},q;z_{\rm s})$ from the EMBS channel may be simply given by (e.g., Cao et al. 2018; Mapelli & Giacobbo 2018; Zhao & Lu 2021)
@@
where $\dot{\psi}(Z;{z^{\prime}})$ is the cosmic star formation rate density (SFR) with metallicity $Z$ at the formation redshift $z^{\prime}$, which can be tracked from cosmological hydro-dynamic-simulations for galaxy formation and evolution, such as Illustris-TNG (Pillepich et al. 2018), EAGLE (Schaye et al. 2015), and GADGET (Springel et al. 2021) or N-body simulations like Millennium (Springel et al. 2005) implemented with semi-analytical galaxy formation models (e.g., Jiang et al. 2019a). The notation $\phi(m_{\ast})$ is the initial mass function (IMF) and $P(q_{\ast})$ is the initial mass ratio of the progenitor binary. The vector function of $g(m_{\ast},q_{\ast},Z)=(m_{1},q)$ describes the mapping between the initial zero-age main sequence binaries and the remnant binaries, which can be obtained from the CBPS simulation. In practice, the evolution of progenitor binaries are simulated by considering several important physical processes, such as mass and angular momentum transfer, common-envelope evolution, natal kick, and mass ejection during SN explosions (e.g., Hurley et al. 2002; Belczynski et al. 2007; Dominik et al. 2012, 2013; Wang et al. 2016; Belczynski et al. 2016; Chu et al. 2022; Riley et al. 2022; Sgalletta et al. 2023; Iorio et al. 2023; Chu et al. 2024; Mandel et al. 2025; Chen et al. 2025b; Wang et al. 2026). Then the cosmic merger rate density of CBCs $R\left(m_{1},q;z_{\rm s}\right)$ can be estimated by Equation (45) with the time delay $\tau_{\rm d}$ determined in the simulation.
The gravitational interactions of (binary) compact stars in dense stellar systems may also lead to the formation of sBBHs (the dynamical channel) and their subsequent mergers due to GW decay, which may contribute significantly to the sBBH mergers, especially for those with large masses (e.g., Sigurdsson & Hernquist 1993; Portegies Zwart & McMillan 2000; Rodriguez et al. 2016; von Zeipel 1910; Kozai 1962; Lidov 1962; Naoz 2016). Such dense stellar systems include globular clusters, open clusters and nuclear clusters. In principle, the sBBH merger rate density due to dynamical channel can be estimated by full dynamical simulations for sBBH formation in globular clusters, which is however computationally expensive. An economic way to estimate the merger rate density is to combine the results of sBBH formation given by numerical simulations with simple models for the formation and evolution of globular clusters and other systems over cosmic time. For example, the merger rate density from the dynamical channel for globular clusters may be roughly estimated as (e.g., Zhao & Lu 2021):
@@
where $\frac{d\dot{M}_{\mathrm{G}}}{d\log M_{\mathrm{H}}}$ is the comoving SFR in globular clusters per galaxies of a given halo mass $M_{\mathrm{H}}$ at given redshift $z(\tau)$ (or a given formation time $\tau$ ). The function $P(M_{\rm G})$ is the cluster initial mass function, $\left\langle M_{\mathrm{G}}\right\rangle$} is the mean initial mass of a globular cluster, while $R(r_{\rm v},M_{\rm G},t)$ is the merger rate of sBBHs in a globular cluster with initial virial radius $r_{\rm v}$ and mass $M_{\rm G}$ at time $t(z_{\rm s})$. For simplicity, the specific form for $\frac{d\dot{M}_{\mathrm{G}}}{d\log M_{\mathrm{H}}}$ and $R(r_{\rm v},M_{\rm G})$ can be assumed as the explicit function given in Rodriguez & Loeb (2018). Here we also comment that the dynamical channel of BNS mergers (Fragione et al. 2019) may be quite inefficient, since NSs are not heavy enough to mostly sink into the centers of globular clusters where the dynamical interactions are most effective (Belczynski et al. 2016). Notably, we did not take into account the case of nuclear clusters, which may also be promising environments for BH-BH dynamical formation due to their deeper potential (e.g., O’Leary et al. 2009; Rasskazov & Kocsis 2019).
The sBBH mergers may be also produced in massive accretion disks in Active-Galactic-Nucleus (AGN) around SMBHs (denoted as AGN channel), which was proposed for the origin of some LVK events, such as GW190521, with extremely large masses that cannot be produced directly from stars due to pair instability supernovae explosion (e.g., McKernan et al. 2012; Bartos et al. 2017; Stone et al. 2017; Zhang et al. 2019; Wang et al. 2021a; Samsing et al. 2022; Gautham Bhaskar et al. 2022; Chen & Dai 2024; Ma et al. 2025). A large number of massive stars can be produced in the AGN disk due to the gravitational instability, which may evolve to BHs and dynamically interact with their gaseous environment and other BHs to form sBBH mergers. Recently, it attracts a lot of interests in this channel as it not only can explain the formation of GW190521-like massive sBBHs but also lead to rich EM signals. However, estimating the merger rate density of sBBHs via this channel is difficult due to their complex environments. However, it is plausible to assume that the merger rate density $R(z_{\rm s})$ via the AGN-MBH channel at redshift $z_{\rm s}$ is proportional to the total rate of accretion of mass onto MBHs at that redshift (Chen et al. 2022b),
@@
where $\Psi(L_{\rm bol},z)$ is the bolometric luminosity function of QSOs, $\epsilon$ is the radiative efficiency of the accretion processes and assumed to be the canonical value $0.1$ (e.g., Yu & Tremaine 2002; Yu & Lu 2004; Marconi et al. 2004; Hopkins et al. 2007; Yu & Lu 2008; Körding et al. 2008; Shankar et al. 2009; Izquierdo-Villalba et al. 2022a).
Note here that there are many complexities within the above procedures, which may introduce large uncertainties. Nevertheless, LVK observations have already put strong constraints on the local merger rate densities, i.e., $19_{-5}^{+7}\rm Gpc^{-3}yr^{-1}$ for sBBHs, $89_{-67}^{+159}\rm Gpc^{-3}yr^{-1}$ for BNSs, and $23_{-13}^{+20}\rm Gpc^{-3}yr^{-1}$ for BHNSs, respectively (The LIGO Scientific Collaboration et al. 2025e). Probably most BNS and BHNS mergers are produced by the EMBS channel, therefore, the local merger rate densities of BNSs and BHNSs may be rescaled by those obtained by LVK observations. For the dynamical channel, however, the value of $R^{D}_{\bullet\bullet}$ is quite difficult to estimate. Nevertheless, Rodriguez & Loeb (2018) obtained $R^{D}_{\bullet\bullet}\sim 14_{-10}^{+4}\rm Gpc^{-3}yr^{-1}$ using a simple model for globular clusters, which may be adopted as a rough calibration. For open clusters and nuclear clusters, it is also possible to estimate the merger rate density with similar implementation in Equation (47), but with substantial uncertainties. For the AGN channel, it is also challenging to estimate the merger rate density accurately due to various complications, though Equation (48) offers a simplified evolutionary approximation. Among these complexities, the most significant source of uncertainty lies in the local merger rate density estimates, which span an exceptionally wide range—from $0.02$ to $60$ Gpc-3 yr-1 (e.g., Stone et al. 2017; McKernan et al. 2018; Fragione et al. 2019; Tagawa et al. 2020; Ishibashi & Gröbner 2024).
Figure 4: The cosmic merger rate density evolution $R(z_{\rm s})$ of compact binary coalescence, including BNS (blue), BHNS (brown), and BBH produced via EMBS (green) and dynamical (red) channel (e.g., Belczynski et al. 2020; Zhao & Lu 2021). The results are rescaled by the latest constraint on the local merger rate density from GWTC-4.0 (The LIGO Scientific Collaboration et al. 2025e), i.e., $R^{\rm E}\sim 19\rm Gpc^{-3}yr^{-1}$ for sBBHs produced via EMBS channel, $89\rm Gpc^{-3}yr^{-1}$ for BNSs, and $23\rm Gpc^{-3}yr^{-1}$ for BHNSs. The merger rate density for BBH produced via dynamical channel are rescaled by $R^{\rm D}\sim 14\rm Gpc^{-3}yr^{-1}$ according to Rodriguez & Loeb (2018).
@@ 3.1.2 Supermassive Binary Black Holes
SMBBHs/MBBHs are natural products of galactic mergers as SMBHs are ubiquitous in galactic centers (e.g., Kormendy & Ho 2013). The population properties and merger rate of SMBBHs/MBBHs are set by the cosmic history of galaxy mergers, the masses (and spins) of the central SMBHs in the progenitor galaxies, and the subsequent orbital evolution of SMBBH pairs during the merger and within merger remnant. To date, due to the absence of robust detections of individual SMBBHs from either EM (e.g., Runnoe et al. 2017; Guo et al. 2019; Chen et al. 2023; Huijse et al. 2025) or PTA (e.g., Agazie et al. 2023b; EPTA Collaboration et al. 2024) observations, the intrinsic population properties and merger rate of SMBBHs cannot yet be inferred from observations directly (see Wang & Li 2020; Lu 2025b). Therefore, estimates of SMBBH population properties and merger rates must rely on theoretical modelling, either semi-analytical approaches that combine observational constraints on the model ingredients (e.g., Sesana 2013b; Chen et al. 2020) or fully hydrodynamic cosmological simulations of galaxy formation and evolution (e.g., Crain et al. 2015; McAlpine et al. 2016; Springel et al. 2018; Pillepich et al. 2018). The former offers great flexibility for exploring large parameter spaces, while the latter captures more detailed galaxy formation physics and the complex interactions between SMBHs/SMBBHs with their hosts. Below we introduce the framework for a semi-analytic SMBBH model based on Chen et al. (2020) with detailed consideration of the SMBBH orbital evolution, while for the SMBBH model through hydrodynamic cosmological simulations one may find in Kelley et al. (2017) and Katz et al. (2020).
@@
The functions $p_{\rm BH}(M_{\rm BH},q_{\rm BH}|M_{*},q_{*},z^{\prime})$ and $p_{f}(f_{\rm GW},\tau|M_{*},q_{*},M_{\rm BH},q_{\rm BH},z^{\prime})$ in Equation (49) are two probability distribution functions to take into account the BH-galaxy relationship, and the SMBBH orbit evolution. The first one is defined so that $p_{\rm BH}(M_{\rm BH},q_{\rm BH}|M_{*},q_{*},z^{\prime})dM_{\rm BH}dq_{\rm BH}$ represents the probability that a host merger at redshift $z^{\prime}$ characterized by $(M_{*},q_{*})$ leads to a SMBBH merger characterized by $(M_{\rm BH},q_{\rm BH})$ with $M_{\rm BH}$ representing the total mass of the SMBBHs, which is inferred by the MBH–host galaxy scaling relation (e.g., Garofalo et al. 2023); The second one is defined so that $p_{f}(f_{\rm GW},\tau|M_{*},q_{*},M_{\rm BH},q_{\rm BH},z^{\prime})df_{\rm GW}$ represents the probability that a host merger characterized by parameters $(M_{*},q_{*},M_{\rm BH},q_{\rm BH})$ at redshift $z^{\prime}$ leads to an SMBBH emitting GW at frequency range $f_{\rm GW}\rightarrow f_{\rm GW}+df_{\rm GW}$ at a time $\tau$ after the host merger. This term encodes the key dynamical processes that SMBBHs experience as they evolve from galaxy pairing to final coalescence. Briefly speaking, the evolution tracks of an SMBBH inside a galaxy merger remnant can be divided into several stages, including (1) dynamical friction stage, when the orbits shrinks mainly due to distant star field and gas particle scattering (e.g., Milosavljević & Merritt 2001; Yu 2002); (2) binary stage with orbital decay governed by interactions with stars and gas material in its environment, first jointly by dynamical friction and three-body scattering of stars passing through the SMBBH (non-hard binary) (e.g., Quinlan 1996; Yu 2002), then dominated by the three-body star scattering (hard binary) (e.g., Quinlan 1996; Yu 2002) and/or by viscous torque from an accretion disk around the binary if the merger occurs in a gas-rich environment; (3) GW radiation dominant stage, with the orbit decay rapidly driving by the GW radiation toward the final coalescence (responsible for nanohertz GW insprialing SMBBHs). The SMBBH orbital evolution have been studied extensively and numerous mechanisms have been proposed to accelerate orbit decay and drive coalescence on timescales shorter than the cosmic (Hubble) time (e.g., Lu 2025b). Assuming the delay between galaxy merger and SMBBH merger is substantially smaller than the Hubble time (e.g., $\sim 10^{9}$ yr or shorter), which may be true for most SMBBH major mergers in small galaxies, then the merger rate of SMBBHs can be estimated as

### 16 May 2026 v2

@@ 2.2.2 Wave Optics
In the wave-optics regime, i.e., ${\lambda_{\rm GW}}/{\lambda_{\rm E}}\gtrsim 1$, the diffraction of the GWs becomes important, which may also lead to changes of the waveform in both amplitude and phase. Here we briefly review the basic theory of the gravitational lensing of GWs in the wave-optics regime, focusing on the typical weak-field case. Extensions to strong-field wave optics can be found in Chan et al. (2025) and Saketh et al. (2026). Figure 2 shows a schematic diagram for the lensing of a GW event by a (dark matter) minihalo in the wave-optics regime. More detailed discussions can be seen in Takahashi and Nakamura (2003).
@@
where $g^{(\rm B)}_{\mu\nu}$ is the background metric tensor. Notably, this expression is valid for static lens, while the case of spinning-lens can be treated by adding an angular-momentum related term as in Prabhu et al. (2025). Consider a linear perturbation $h_{\mu\nu}$ in the background metric $g^{(\rm B)}_{\mu\nu}$, the metric is then
@@
where $n_{i}=0$, $0.5$, and $1$ when $x_{j}$ is the minimum (Type I), saddle (Type II), and maximum (Type III) points of $t_{\rm d}(\boldsymbol{x,y})$, respectively, representing the parity of the images. Note that Type-II images acquire an additional phase shift of $\pi/2$ relative to Type-I images, as encoded by the Morse phase term $n_{i}=0.5$. Such a phase shift may introduce characteristic distortions in the observed waveform, especially for signals with more complex morphology, such as higher harmonics, precession, or eccentricity (e.g., Ezquiaga et al. 2021; Wang et al. 2021c; Vijaykumar et al. 2023). This equation reduces to the geometric-optics expression in the short-wavelength limit, implying that geometric-optics is a valid approximation when diffraction can be ignored.
We note here that calculating lensed waveforms in the wave-optics regime via the amplification factor $F(f)$ via Equation (41) presents two key challenges. First, from a theoretical standpoint, Equation (41) is derived using Kirchhoff’s diffraction theory, which inherently assumes that the incident GW is obstructed by the lens and only accounts for scattered wave contribution (see also Fig. 2 for a schematic illustration). When considering that GWs propagate past BHs, fully numerical relativity simulations are necessary to obtain more robust lensed GW waveforms that incorporate the contributions from incident GWs, rather than only scattered components (e.g., He 2021; He and Wu 2022; Yin and He 2024). Second, from a numerical computational perspective, calculating $F(f)$ via Equation (41) in the wave-optics regime is both challenging and computationally expensive owing to highly oscillatory nature of the diffraction integral. Efficient computation of this integral is feasible for axis-symmetric lens profiles (Guo and Lu 2020). For more general lensing profiles, the calculation of the diffraction integral can be done by numerical codes implementing fast Fourier transformation, such as FIONA [^6] https://github.com/ninoephremidze/FIONA (Ephremidze et al. 2026) and GLoW (Villarrubia-Rojo et al. 2025) [^7] https://github.com/miguelzuma/GLoW_public.
@@ 3.1.1 Compact Binary Coalescence
Binary neutron star (BNSs) mergers, black hole-neutron star (BHNS) mergers and a significant fraction of stellar binary black holes (sBBHs) are believed to originate from the evolution of massive binary stars (hereafter, EMBS channel). One may adopt the cosmic merger rate densities directly obtained from GW observations, however, the uncertainties in these measurements are still large due to small number of detected GW events and lack of high redshift GW events. A widely adopted approach is to combine compact binary population synthesis (CBPS) model and cosmological simulation of galaxy formation and evolution (e.g., Dominik et al. 2012, 2013, 2015; de Mink and Belczynski 2015; Mapelli and Giacobbo 2018; Han et al. 2020; Chu et al. 2022; Fragos et al. 2023; Chu et al. 2022; Chen et al. 2024c, 2025c, also see recent reviews by Mandel and Broekgaarden 2022 and Lu 2025a). In principle, the cosmic merger rate density $R^{\rm S}(m_{1},q;z_{\rm s})$ from the EMBS channel may be simply given by (e.g., Cao et al. 2018; Mapelli and Giacobbo 2018; Zhao and Lu 2021)
@@
where $\dot{\psi}(Z;{z^{\prime}})$ is the cosmic star formation rate density (SFR) with metallicity $Z$ at the formation redshift $z^{\prime}$, which can be tracked from cosmological hydro-dynamic-simulations for galaxy formation and evolution, such as Illustris-TNG (Pillepich et al. 2018), EAGLE (Schaye et al. 2015), and GADGET (Springel et al. 2021) or N-body simulations like Millennium (Springel et al. 2005) implemented with semi-analytical galaxy formation models (e.g., Jiang et al. 2019a). The notation $\phi(m_{\ast})$ is the initial mass function (IMF) and $P(q_{\ast})$ is the initial mass ratio of the progenitor binary. The vector function of $g(m_{\ast},q_{\ast},Z)=(m_{1},q)$ describes the mapping between the initial zero-age main sequence binaries and the remnant binaries, which can be obtained from the CBPS simulation. In practice, the evolution of progenitor binaries are simulated by considering several important physical processes, such as mass and angular momentum transfer, common-envelope evolution, natal kick, and mass ejection during SN explosions (e.g., Hurley et al. 2002; Belczynski et al. 2007; Dominik et al. 2012, 2013; Wang et al. 2016; Belczynski et al. 2016; Chu et al. 2022; Riley et al. 2022; Sgalletta et al. 2023; Iorio et al. 2023; Chu et al. 2024; Mandel et al. 2025; Chen et al. 2025c; Wang et al. 2026). Then the cosmic merger rate density of CBCs $R\left(m_{1},q;z_{\rm s}\right)$ can be estimated by Equation (45) with the time delay $\tau_{\rm d}$ determined in the simulation.
The gravitational interactions of (binary) compact stars in dense stellar systems may also lead to the formation of sBBHs (the dynamical channel) and their subsequent mergers due to GW decay, which may contribute significantly to the sBBH mergers, especially for those with large masses (e.g., Sigurdsson and Hernquist 1993; Portegies Zwart and McMillan 2000; Rodriguez et al. 2016; von Zeipel 1910; Kozai 1962; Lidov 1962; Naoz 2016). Such dense stellar systems include globular clusters, open clusters and nuclear clusters. In principle, the sBBH merger rate density due to dynamical channel can be estimated by full dynamical simulations for sBBH formation in globular clusters, which is however computationally expensive. An economic way to estimate the merger rate density is to combine the results of sBBH formation given by numerical simulations with simple models for the formation and evolution of globular clusters and other systems over cosmic time. For example, the merger rate density from the dynamical channel for globular clusters may be roughly estimated as (e.g., Zhao and Lu 2021):
@@
where $\frac{d\dot{M}_{\mathrm{G}}}{d\log M_{\mathrm{H}}}$ is the comoving SFR in globular clusters per galaxies of a given halo mass $M_{\mathrm{H}}$ at given redshift $z(\tau)$ (or a given formation time $\tau$ ). The function $P(M_{\rm G})$ is the cluster initial mass function, $\left\langle M_{\mathrm{G}}\right\rangle$} is the mean initial mass of a globular cluster, while $R(r_{\rm v},M_{\rm G},t)$ is the merger rate of sBBHs in a globular cluster with initial virial radius $r_{\rm v}$ and mass $M_{\rm G}$ at time $t(z_{\rm s})$. For simplicity, the specific form for $\frac{d\dot{M}_{\mathrm{G}}}{d\log M_{\mathrm{H}}}$ and $R(r_{\rm v},M_{\rm G})$ can be assumed as the explicit function given in Rodriguez and Loeb (2018). Here we also comment that the dynamical channel of BNS mergers (Fragione et al. 2019) may be quite inefficient, since NSs are not heavy enough to mostly sink into the centers of globular clusters where the dynamical interactions are most effective (Belczynski et al. 2016). Notably, we did not take into account the case of nuclear clusters, which may also be promising environments for BH-BH dynamical formation due to their deeper potential (e.g., O’Leary et al. 2009; Rasskazov and Kocsis 2019).
The sBBH mergers may be also produced in massive accretion disks in Active-Galactic-Nucleus (AGN) around SMBHs (denoted as AGN channel), which was proposed for the origin of some LVK events, such as GW190521, with extremely large masses that cannot be produced directly from stars due to pair instability supernovae explosion (e.g., McKernan et al. 2012; Bartos et al. 2017; Stone et al. 2017; Zhang et al. 2019; Wang et al. 2021a; Samsing et al. 2022; Gautham Bhaskar et al. 2022; Chen and Dai 2024; Ma et al. 2025). A large number of massive stars can be produced in the AGN disk due to the gravitational instability, which may evolve to BHs and dynamically interact with their gaseous environment and other BHs to form sBBH mergers. Recently, it attracts a lot of interests in this channel as it not only can explain the formation of GW190521-like massive sBBHs but also lead to rich EM signals. However, estimating the merger rate density of sBBHs via this channel is difficult due to their complex environments. However, it is plausible to assume that the merger rate density $R(z_{\rm s})$ via the AGN-MBH channel at redshift $z_{\rm s}$ is proportional to the total rate of accretion of mass onto MBHs at that redshift (Chen et al. 2022b),
@@
where $\Psi(L_{\rm bol},z)$ is the bolometric luminosity function of QSOs, $\epsilon$ is the radiative efficiency of the accretion processes and assumed to be the canonical value $0.1$ (e.g., Yu and Tremaine 2002; Yu and Lu 2004; Marconi et al. 2004; Hopkins et al. 2007; Yu and Lu 2008; Körding et al. 2008; Shankar et al. 2009; Izquierdo-Villalba et al. 2022a).
Note here that there are many complexities within the above procedures, which may introduce large uncertainties. Nevertheless, LVK observations have already put strong constraints on the local merger rate densities, i.e., $19_{-5}^{+7}\rm Gpc^{-3}yr^{-1}$ for sBBHs, $89_{-67}^{+159}\rm Gpc^{-3}yr^{-1}$ for BNSs, and $23_{-13}^{+20}\rm Gpc^{-3}yr^{-1}$ for BHNSs, respectively (The LIGO Scientific Collaboration et al. 2025b). Probably most BNS and BHNS mergers are produced by the EMBS channel, therefore, the local merger rate densities of BNSs and BHNSs may be rescaled by those obtained by LVK observations. For the dynamical channel, however, the value of $R^{D}_{\bullet\bullet}$ is quite difficult to estimate. Nevertheless, Rodriguez and Loeb (2018) obtained $R^{D}_{\bullet\bullet}\sim 14_{-10}^{+4}\rm Gpc^{-3}yr^{-1}$ using a simple model for globular clusters, which may be adopted as a rough calibration. For open clusters and nuclear clusters, it is also possible to estimate the merger rate density with similar implementation in Equation (47), but with substantial uncertainties. For the AGN channel, it is also challenging to estimate the merger rate density accurately due to various complications, though Equation (48) offers a simplified evolutionary approximation. Among these complexities, the most significant source of uncertainty lies in the local merger rate density estimates, which span an exceptionally wide range—from $0.02$ to $60$ Gpc-3 yr-1 (e.g., Stone et al. 2017; McKernan et al. 2018; Fragione et al. 2019; Tagawa et al. 2020; Ishibashi and Gröbner 2024).
Figure 4: The cosmic merger rate density evolution $R(z_{\rm s})$ of compact binary coalescence, including BNS (blue), BHNS (brown), and BBH produced via EMBS (green) and dynamical (red) channel (e.g., Belczynski et al. 2020; Zhao and Lu 2021). The results are rescaled by the latest constraint on the local merger rate density from GWTC-4.0 (The LIGO Scientific Collaboration et al. 2025b), i.e., $R^{\rm E}\sim 19\rm Gpc^{-3}yr^{-1}$ for sBBHs produced via EMBS channel, $89\rm Gpc^{-3}yr^{-1}$ for BNSs, and $23\rm Gpc^{-3}yr^{-1}$ for BHNSs. The merger rate density for BBH produced via dynamical channel are rescaled by $R^{\rm D}\sim 14\rm Gpc^{-3}yr^{-1}$ according to Rodriguez and Loeb (2018).
@@ 3.1.2 Supermassive Binary Black Holes
SMBBHs/MBBHs are natural products of galactic mergers as SMBHs are ubiquitous in galactic centers (e.g., Kormendy and Ho 2013). The population properties and merger rate of SMBBHs/MBBHs are set by the cosmic history of galaxy mergers, the masses (and spins) of the central SMBHs in the progenitor galaxies, and the subsequent orbital evolution of SMBBH pairs during the merger and within merger remnant. To date, due to the absence of robust detections of individual SMBBHs from either EM (e.g., Runnoe et al. 2017; Guo et al. 2019; Chen et al. 2023a; Huijse et al. 2025) or PTA (e.g., Agazie et al. 2023b; EPTA Collaboration et al. 2024) observations, the intrinsic population properties and merger rate of SMBBHs cannot yet be inferred from observations directly (see Wang and Li 2020; Lu 2025b). Therefore, estimates of SMBBH population properties and merger rates must rely on theoretical modelling, either semi-analytical approaches that combine observational constraints on the model ingredients (e.g., Sesana 2013b; Chen et al. 2020) or fully hydrodynamic cosmological simulations of galaxy formation and evolution (e.g., Crain et al. 2015; McAlpine et al. 2016; Springel et al. 2018; Pillepich et al. 2018). The former offers great flexibility for exploring large parameter spaces, while the latter captures more detailed galaxy formation physics and the complex interactions between SMBHs/SMBBHs with their hosts. Below we introduce the framework for a semi-analytic SMBBH model based on Chen et al. (2020) with detailed consideration of the SMBBH orbital evolution, while for the SMBBH model through hydrodynamic cosmological simulations one may find in Kelley et al. (2017) and Katz et al. (2020).
@@
The functions $p_{\rm BH}(M_{\rm BH},q_{\rm BH}|M_{*},q_{*},z^{\prime})$ and $p_{f}(f_{\rm GW},\tau|M_{*},q_{*},M_{\rm BH},q_{\rm BH},z^{\prime})$ in Equation (49) are two probability distribution functions to take into account the BH-galaxy relationship, and the SMBBH orbit evolution. The first one is defined so that $p_{\rm BH}(M_{\rm BH},q_{\rm BH}|M_{*},q_{*},z^{\prime})dM_{\rm BH}dq_{\rm BH}$ represents the probability that a host merger at redshift $z^{\prime}$ characterized by $(M_{*},q_{*})$ leads to a SMBBH merger characterized by $(M_{\rm BH},q_{\rm BH})$ with $M_{\rm BH}$ representing the total mass of the SMBBHs, which is inferred by the MBH–host galaxy scaling relation (e.g., Garofalo et al. 2023); The second one is defined so that $p_{f}(f_{\rm GW},\tau|M_{*},q_{*},M_{\rm BH},q_{\rm BH},z^{\prime})df_{\rm GW}$ represents the probability that a host merger characterized by parameters $(M_{*},q_{*},M_{\rm BH},q_{\rm BH})$ at redshift $z^{\prime}$ leads to an SMBBH emitting GW at frequency range $f_{\rm GW}\rightarrow f_{\rm GW}+df_{\rm GW}$ at a time $\tau$ after the host merger. This term encodes the key dynamical processes that SMBBHs experience as they evolve from galaxy pairing to final coalescence. Briefly speaking, the evolution tracks of an SMBBH inside a galaxy merger remnant can be divided into several stages, including (1) dynamical friction stage, when the orbits shrinks mainly due to distant star field and gas particle scattering (e.g., Milosavljević and Merritt 2001; Yu 2002); (2) binary stage with orbital decay governed by interactions with stars and gas material in its environment, first jointly by dynamical friction and three-body scattering of stars passing through the SMBBH (non-hard binary) (e.g., Quinlan 1996; Yu 2002), then dominated by the three-body star scattering (hard binary) (e.g., Quinlan 1996; Yu 2002) and/or by viscous torque from an accretion disk around the binary if the merger occurs in a gas-rich environment; (3) GW radiation dominant stage, with the orbit decay rapidly driving by the GW radiation toward the final coalescence (responsible for nanohertz GW insprialing SMBBHs). The SMBBH orbital evolution have been studied extensively and numerous mechanisms have been proposed to accelerate orbit decay and drive coalescence on timescales shorter than the cosmic (Hubble) time (e.g., Lu 2025b). Assuming the delay between galaxy merger and SMBBH merger is substantially smaller than the Hubble time (e.g., $\sim 10^{9}$ yr or shorter), which may be true for most SMBBH major mergers in small galaxies, then the merger rate of SMBBHs can be estimated as